Find the two-digit number whose tens digit is 3 less than the units digit. The original number is 6 more than 4 times the sum of the digits.
Answer by mathmaven53(29)

Let x be units digit then by condition of problem tens digit is x-3
The origonal number is 10(x-3) + x
The sum of the digits is x-3 + x
The origonal number is six more than four times the sum of the digits
10(x-3)+x = 4(x-3 + x) + 6
Simplify the left and right sides
11x - 30 = 8x - 6
From this
3x = 24
x = 8

The number is 58